A country is described by the Solow model, with a production function of y = k1/2.
Suppose that k is equal to 225. The fraction of output invested is 60%. The depreciation
rate is 5%. Is the country at its steady-state level of output per worker, above the
steady state, or below the steady state? Explain how you reached your conclusion
1
Expert's answer
2016-02-10T00:01:18-0500
A country is described by the Solow model. If a production function of y = k1/2 and k is equal to 225, then y = 225^1/2 = 15. If the fraction of output invested is 60%, then the investment is 15*0.6 = 9. The depreciation
rate is 5%. The country at its steady-state level of output per worker, because the economy is of relatively stable size, has stable population and stable consumption that remain at or below carrying capacity.